# Main-belt and Trojan Asteroid Phase Curves from the ATLAS Survey

### v1

@@ Abstract
pubyear: 2023pagerange: Main-belt and Trojan Asteroid Phase Curves from the ATLAS SurveyC
@@ 2.2 The rockAtlas database
For this study we generated a database of ATLAS detections matched to known Solar System minor planets. The rockAtlas database (Young, 2023) has been processing all ATLAS observations for a fixed list of known asteroids since 2017. The initial list of asteroids to be tracked was $\sim 445,000$ asteroids from the astorb  [^2] https://asteroid.lowell.edu/main/astorb/ dataset (Moskovitz et al., 2022) with well defined orbits at that time. Accurate orbital parameters were required as rockAtlas calculates the ephemerides of these objects and matches them to ATLAS observations. When rockAtlas is running, a night’s worth of observations is downloaded from the ATLAS automated detection and photometry pipeline - dophot (Tonry et al., 2018). dophot provides details on all ATLAS detections that night, such as time, position, magnitude etc. These detections consist of moving Solar System bodies, static transients and false detections. rockAtlas calculates precise ephemerides for all asteroids in its watch list at the mid-time of each exposure using pyephem  [^3] https://rhodesmill.org/pyephem/ and orbfit  [^4] http://adams.dm.unipi.it/orbfit/. These ephemerides are then cross-matched to the dophot detection positions using a HEALpix grid (Gorski et al., 2005; Zonca et al., 2019). A dophot detection is associated with a moving object, provided the separation from the predicted ephemeris at the exposure mid-time is less than $5\,$\mathrm{a}\mathrm{r}\mathrm{c}\mathrm{s}\mathrm{e}\mathrm{c}$$, and it is added to the rockAtlas database. Together with the ephemerides, pyephem and orbfit provide the viewing geometry for each observation including heliocentric and geocentric distances, the Sun-asteroid-Earth phase angle and galactic latitude, which are appended to the database.
@@
Figure 1: Two dimensional histogram showing the number distribution of the maximum observed phase angle of asteroids in the rockAtlas database ($o$-filter) as a function of the object’s semimajor axis (left-hand panel). The maximum phase angle for a zero-inclination circular orbit, $\text{max}(\alpha)=\sin^{-1}{1/a}$, is shown as a dotted red line. The total distribution of maximum phase angle for the whole dataset is shown as a vertical histogram in the right-hand panel. The Main Asteroid Belt dominates the database; these asteroids are typically observed up to a maximum phase angle of $\gtrsim 20^{\circ}$.
@@ 2.4 Phase curve fitting algorithm
We then performed additional filtering of outlying data to each individual apparition, considering the $o$ and $c$-filters separately. To each apparition we fitted the B89 $HG$ phase curve model using a Levenberg-Marquadt non-linear least squares algorithm (astropy; Astropy Collaboration et al., 2022). For this first stage of phase curve fitting the data points were not weighted by their photometric uncertainties, in order to reduce the influence of extreme outliers in brightness with very low uncertainties. The initial estimate for $H$ was taken from astorb and colour corrected from $V$-band to the ATLAS filter of the observations. This conversion depends on an asteroid’s colour and taxonomic class but we use the values from Heinze et al. (2021) which are derived from the mean colour of S and C type asteroids.
@@ 2.6 Apparition effects, absolute magnitudes and asteroid colours in ATLAS filters
Figure 2: ATLAS phase curves of asteroid (766) Moguntia, an object displaying a large apparition effect. Observations in the $o$-filter are shown in the left-hand panel. Each apparition has been colour coded and outlying photometry/rejected apparitions are shown as small black markers (some of which are obscured by the figure legend). The limits of the $y$-axis have been set to focus on retained photometry, ignoring any more extreme outliers. The corresponding $HG$ phase curve fit for each apparition is shown by the matching coloured curves, with the primary apparition indicated by the coloured star. The median $H$ of the retained apparitions is indicated by the dashed black curve with parameters: $H_{o}=9.68$, $G=0.21$. Observations in the $c$-filter are shown in the right-hand panel, where marker/line colour have the same meanings as the left panel. The large shifts in $H$ between apparitions are clear ($F_{\text{app}}=5.9$; equation 8).
Figure 3: Phase curves for (4986) Osipovia, a typical MBA observed by ATLAS. As in figure 2 the left and right-hand panels show the $o$ and $c$-filter observations respectively. Marker/line colour indicates the apparition to which each observation belongs, with the primary apparition marker by the coloured star and the median value of $H$ indicated by the black dashed curve with parameters: $H_{o}=13.21$, $G=0.30$
@@ 3.1 ATLAS phase curve database
Moguntia 766 1029 363 59432.3 5 21.24 21.22 19.68 19.73 3.02 0.10 10.090543
Priamus 884 1297 420 59523.2 6 12.60 12.59 11.53 12.26 5.20 0.13 8.906102
Marietta 2144 1309 433 59600.2 5 21.55 21.36 21.36 20.44 2.87 0.06 2.833143
Yatskiv 2728 1192 335 59600.2 5 29.29 29.21 28.32 28.02 2.46 0.17 2.610173
Osipovia 4986 652 246 59596.5 5 27.46 27.37 25.76 24.92 2.56 0.21 5.611378
1991 EN 9857 1065 314 59598.3 6 11.57 11.45 10.68 9.28 5.16 0.03 19.612052
@@
Table 1: Overview of the observations made by ATLAS and collected by rockAtlas, for a selected sample of 6 asteroids (this table will be available in its entirety online). We provide the name and MPC number of each asteroid, alongside the number of detections in the $o$ and $c$-filters ($N_{o}$, $N_{c}$) and the date of the last observation used in this work, $\text{max}(t)$. The maximum phase angle observed in each filter is given by $\text{max}(\alpha_{o})$ and $\text{max}(\alpha_{c})$, and the maximum minus minimum phase angle range for each filter is $\Delta\alpha_{o}$, $\Delta\alpha_{c}$. Finally, the orbital elements obtained from astorb are also included; semimajor axis $a$, eccentricity $e$ and inclination $i$.
Name
Number
$n_{\textrm{app}}$
$t_{\textrm{app},0}$ (mjd)
max($\alpha_{\textrm{app},o}$)
max($\alpha_{\textrm{app},c}$)
$N_{\textrm{fit},o}$
$N_{\textrm{fit},c}$
$H_{o}$
$H_{c}$
$\sigma_{Ho}$
$\sigma_{Hc}$
$G$
$\sigma_{G}$
Primary Apparition
[…]
Moguntia
766
0
57380.7
17.90
18.00
47
59
9.28
9.66
6.93e-03
3.65e-03
0.21
NaN
0
1
1
Moguntia
766
1
57883.6
18.09
17.34
146
20
10.09
10.67
8.86e-03
2.34e-02
0.21
NaN
0
1
0
Moguntia
766
2
58298.6
20.52
20.53
199
49
9.25
9.61
2.00e-03
4.54e-03
0.21
NaN
0
0
0
Moguntia
766
3
58733.6
21.24
21.22
232
70
10.15
10.51
8.96e-03
1.61e-02
0.21
NaN
0
1
1
Moguntia
766
4
59190.6
18.29
18.46
244
138
9.27
9.63
2.42e-03
1.22e-03
0.21
2.87e-03
1
1
1
Priamus
884
0
57384.4
9.92
9.94
41
74
8.59
8.99
1.93e-02
1.54e-02
0.23
NaN
0
0
0
Priamus
884
1
57673.6
10.30
10.30
145
88
8.43
8.86
5.70e-03
1.13e-02
0.23
NaN
0
1
1
Priamus
884
2
58037.6
10.94
10.78
179
22
8.39
8.72
3.90e-03
1.14e-02
0.23
NaN
0
0
0
Priamus
884
3
58460.6
11.82
11.82
292
35
8.39
8.75
9.36e-03
9.19e-03
0.23
1.48e-02
1
1
1
Priamus
884
4
58867.6
12.60
12.59
200
80
8.50
8.86
7.11e-03
1.27e-02
0.23
NaN
0
1
1
Priamus
884
5
59255.6
12.05
12.37
20
14
8.55
8.96
2.54e-02
1.50e-02
0.23
NaN
0
0
0
Marietta
2144
0
57481.6
20.26
20.39
45
112
10.99
11.39
2.12e-02
1.33e-02
0.21
NaN
0
1
1
Marietta
2144
1
57943.6
21.55
21.36
175
27
11.06
11.39
1.14e-02
2.48e-02
0.21
NaN
0
1
0
Marietta
2144
2
58417.6
20.14
20.28
256
52
10.96
11.42
8.33e-03
1.77e-02
0.21
NaN
0
1
1
Marietta
2144
3
58846.6
19.45
19.45
309
126
11.08
11.41
8.92e-03
1.27e-02
0.21
NaN
0
1
1
Marietta
2144
4
59300.6
20.58
20.47
410
90
11.02
11.41
1.28e-02
1.39e-02
0.21
1.58e-02
1
1
0
Yatskiv
2728
0
57400.5
22.19
21.77
49
75
12.42
12.56
1.66e-02
2.09e-02
0.11
NaN
0
1
1
Yatskiv
2728
1
57826.6
28.61
27.77
113
13
12.29
12.59
1.07e-02
1.65e-02
0.11
NaN
0
1
0
Yatskiv
2728
2
58385.6
20.64
19.43
131
28
12.30
12.65
1.23e-02
3.43e-02
0.11
NaN
0
1
0
Yatskiv
2728
3
58790.6
23.23
23.01
231
66
12.50
12.63
1.24e-02
1.82e-02
0.11
NaN
0
1
1
Yatskiv
2728
4
59276.6
29.29
29.21
404
85
12.31
12.56
1.55e-02
1.09e-02
0.11
1.35e-02
1
1
1
Osipovia
4986
0
57591.6
27.46
27.37
48
46
13.21
13.63
1.53e-02
1.65e-02
0.30
NaN
0
1
1
Osipovia
4986
1
58037.6
24.71
23.86
158
37
13.23
13.63
1.01e-02
2.07e-02
0.30
NaN
0
1
1
Osipovia
4986
2
58515.7
NaN
NaN
NaN
NaN
NaN
NaN
NaN
NaN
NaN
NaN
NaN
0
0
Osipovia
4986
3
58974.6
25.95
25.79
191
96
13.19
13.58
1.82e-02
9.53e-03
0.30
2.31e-02
1
1
1
Osipovia
4986
4
59493.6
26.30
24.83
110
35
13.22
13.60
1.10e-02
2.15e-02
0.30
NaN
0
1
0
1991 EN
9857
0
57461.6
NaN
NaN
NaN
NaN
NaN
NaN
NaN
NaN
NaN
NaN
NaN
0
0
1991 EN
9857
1
57882.6
9.72
8.05
36
7
10.04
10.48
2.03e-02
6.54e-02
0.16
NaN
0
0
0
1991 EN
9857
2
58269.5
11.38
11.37
140
45
10.16
10.49
1.03e-02
1.41e-02
0.16
NaN
0
1
1
1991 EN
9857
3
58613.6
11.57
11.45
255
66
10.09
10.40
4.74e-03
1.36e-02
0.16
NaN
0
1
1
1991 EN
9857
4
58997.6
11.40
11.39
209
109
10.09
10.44
3.96e-03
7.08e-03
0.16
NaN
0
1
1
1991 EN
9857
5
59414.6
11.17
10.99
352
46
10.10
10.46
1.36e-02
1.34e-02
0.16
2.14e-02
1
1
1
[…]
@@
Figure 4: Log-scale distributions of $H$ (left), $G$ (middle) and $H_{c}-H_{o}$ colour (right) for ATLAS phase curve fits that have passed the quality control checks (section 2.5) and are included in the SAP dataset (section 3.1). $H$ and $G$ are given in the $o$-filter and B89 model. In both figures we show the underlying distributions of S and C type asteroids within the ATLAS dataset (taxonomy from SsODNet; Berthier et al., 2023). The median values of the S and C-type distributions are indicated by the vertical line of matching colour and their values are as follows: $H_{o}(\mathrm{S})=14.83$, $H_{o}(\mathrm{C})=14.52$, $G(\mathrm{S})=0.29$, $G(\mathrm{C})=0.21$, $H_{c}-H_{o}(\mathrm{S})=0.39$, $H_{c}-H_{o}(\mathrm{C})=0.27$. In the middle panel we indicate the boundaries on $G$ for a monotonically increasing B89 phase curve with vertical red dashed lines.
@@
Figure 6: Two dimensional histogram of $H_{c}-H_{o}$ colour vs. $G$. The bimodal distribution of matched S and C-type asteroids (SsODNet; Berthier et al., 2023) is highlighted by orange and green contours respectively, which indicates the 25% level with respect to the largest histogram bin value of their sub-distributions in $H_{c}-H_{o}$ vs. $G$ parameter space. The median values of the S and C-types with ATLAS colours are $G=0.29,0.21$ and $H_{c}-H_{o}=0.39,0.27\,$\mathrm{m}\mathrm{a}\mathrm{g}$$ respectively.
We have also derived the phase curve parameters of asteroids using the P16 $H_{12}G_{12}$ system, where we use the notation $H_{12}$ to denote absolute magnitudes in this system. As described in section 2.5 the quality control cuts were developed on the $HG$ dataset and then applied separately to the residuals of the $H_{12}G_{12}$ phase curve fits. Therefore not all asteroids are present in both datasets. Regardless, both phase curve models show broadly similar results. As before the phase parameter and colour distributions are both bimodal due to the dominant S and C-type asteroids, the former shown in figure 7. The quality control metrics developed for $HG$ do not lead to as strong restrictions on $G_{12}$ for “physical” values between 0 and 1. This appears to be mostly due to inherent differences between the $HG$ and $H_{12}G_{12}$ models. Figure 8 shows that there is a strong correlation between $G_{12}$ and $G$ for objects in our dataset with a phase curve fit in both models. We have verified that this is caused by the different formulations of the two models and is not due to systematic effects in our phase curve fitting procedure. As detailed in appendix B (figure 17), we generated a range of ideal model $HG$ phase curves with $0\leq G\ \leq 1$ and then fitted those simulated observations with the $H_{12}G_{12}$ model. The locus of best fit $G_{12}$ values for phase curves with parameter $G$ is shown in the top panel of figure 8. It can be seen that fitting P16 to observed phase curves naturally gives rise to a range of $G_{12}$ larger than $(0,1)$
Figure 7: Log-scale distribution of P16 $G_{12}$ phase parameters for the quality controlled SAP dataset. Distributions of objects with S and C-type taxonomic classifications from Berthier et al. (2023) are also shown. Similar to figure 4 the median values of the S and C-type distributions are indicated by coloured vertical lines (with values of $G_{12}(\mathrm{S})=0.22$, $G_{12}(\mathrm{C})=0.51$) and the limits $10<G_{12}<$ 1 for a “physical” phase curve are shown as vertical red dashed lines.
@@
Figure 8: Two dimensional histogram showing the relation between the $G_{12}$ and $G$ parameters for objects in our database (upper panel). A 3rd-order polynomial is fitted to the data as shown by the black curve (equation 9a). The red markers indicate the values of $G_{12}$ when a $H_{12}G_{12}$ model is fitted to a $HG$ model (appendix B). This empirical test follows the trend in the ATLAS data closely. The bottom panel shows the two dimensional histogram of the difference in fitted absolute magnitude $H_{12}-H$ vs. $G$. Again, the black curve is a 3rd-order polynomial fit (equation 9b) and the red markers indicate the expected difference in absolute magnitude when $H_{12}G_{12}$ is fitted to $HG$. The horizontal black line indicates the zero value of the $H_{12}-H$ axis.
Figure 9: Two dimensional histogram which shows the variation of $\Delta H=H_{12}-H$ with $H$. The median difference in absolute magnitude between the two phase curve models is $-0.11$ (indicated by the horizontal red line). There is a negative correlation between $\Delta H$ and $H$, as well as a trend for a larger dispersion in $\Delta H$ for fainter objects (see also figure 8). The horizontal black line indicates the zero value of the $H_{12}-H$ axis.

### v2

@@ Abstract
pubyear: 2024pagerange: Main-belt and Trojan Asteroid Phase Curves from the ATLAS SurveyC
@@ 2.2 The rockAtlas database
For this study we generated a database of ATLAS detections matched to known Solar System minor planets. The rockAtlas database (Young, 2018) has been processing all ATLAS observations for a fixed list of known asteroids since 2017. The initial list of asteroids to be tracked was $\sim 445,000$ asteroids from the astorb  [^2] https://asteroid.lowell.edu/main/astorb/ dataset (Moskovitz et al., 2022) with well defined orbits at that time. Accurate orbital parameters were required as rockAtlas calculates the ephemerides of these objects and matches them to ATLAS observations. When rockAtlas is running, a night’s worth of observations is downloaded from the ATLAS automated detection and photometry pipeline - dophot (Tonry et al., 2018). dophot provides details on all ATLAS detections that night, such as time, position, magnitude etc. These detections consist of moving Solar System bodies, static transients and false detections. rockAtlas calculates precise ephemerides for all asteroids in its watch list at the mid-time of each exposure using pyephem  [^3] https://rhodesmill.org/pyephem/ and orbfit  [^4] http://adams.dm.unipi.it/orbfit/. These ephemerides are then cross-matched to the dophot detection positions using a HEALpix grid (Gorski et al., 2005; Zonca et al., 2019). A dophot detection is associated with a moving object, provided the separation from the predicted ephemeris at the exposure mid-time is less than $5\,$\mathrm{a}\mathrm{r}\mathrm{c}\mathrm{s}\mathrm{e}\mathrm{c}$$, and it is added to the rockAtlas database. Together with the ephemerides, pyephem and orbfit provide the viewing geometry for each observation including heliocentric and geocentric distances, the Sun-asteroid-Earth phase angle and galactic latitude, which are appended to the database.
@@
Figure 1: Two dimensional histogram showing the number distribution of the maximum observed phase angle of asteroids in the rockAtlas database ($o$-filter) as a function of the object’s semimajor axis (left-hand panel). The maximum phase angle for a zero-inclination circular orbit, $\text{max}(\alpha)=\sin^{-1}{1/a}$, is shown as a dotted line. The total distribution of maximum phase angle for the whole dataset is shown as a vertical histogram in the right-hand panel. The Main Asteroid Belt dominates the database; these asteroids are typically observed up to a maximum phase angle of $\gtrsim 20^{\circ}$.
@@ 2.4 Phase curve fitting algorithm
We then performed additional filtering of outlying data to each individual apparition, considering the $o$ and $c$-filters separately. To each apparition we fitted the B89 $HG$ phase curve model using a Levenberg-Marquardt non-linear least squares algorithm (astropy; Astropy Collaboration et al., 2022). For this first stage of phase curve fitting the data points were not weighted by their photometric uncertainties, in order to reduce the influence of extreme outliers in brightness with very low uncertainties. The initial estimate for $H$ was taken from astorb and colour corrected from $V$-band to the ATLAS filter of the observations. This conversion depends on an asteroid’s colour and taxonomic class but we use the values from Heinze et al. (2021) which are derived from the mean colour of S and C type asteroids.
@@ 2.6 Apparition effects, absolute magnitudes and asteroid colours in ATLAS filters
Figure 2: ATLAS phase curves of asteroid (766) Moguntia, an object displaying a large apparition effect. Observations in the $o$-filter are shown in the left-hand panel. Each apparition has been colour coded and outlying photometry/rejected apparitions are shown as small markers (some of which are obscured by the figure legend). The limits of the $y$-axis have been set to focus on retained photometry, ignoring any more extreme outliers. The corresponding $HG$ phase curve fit for each apparition is shown by the matching coloured curves, with the primary apparition indicated by the coloured star. The median $H$ of the retained apparitions is indicated by the dashed curve with parameters: $H_{o}=9.68$, $G=0.21$. Observations in the $c$-filter are shown in the right-hand panel, where marker/line colour have the same meanings as the left panel. The large shifts in $H$ between apparitions are clear ($F_{\text{app}}=5.9$; equation 8).
Figure 3: Phase curves for (4986) Osipovia, a typical MBA observed by ATLAS. As in figure 2 the left and right-hand panels show the $o$ and $c$-filter observations respectively. Marker/line colour indicates the apparition to which each observation belongs, with the primary apparition marker by the coloured star and the median value of $H$ indicated by the dashed curve with parameters: $H_{o}=13.21$, $G=0.30$
@@ 3.1 ATLAS phase curve database
Moguntia 766 1029 363 59432.262037 5 21.24 21.22 19.68 19.73 3.02 0.10 10.090543
Priamus 884 1297 420 59523.217303 6 12.60 12.59 11.53 12.26 5.20 0.13 8.906102
Marietta 2144 1309 433 59600.229942 5 21.55 21.36 21.36 20.44 2.87 0.06 2.833143
Yatskiv 2728 1192 335 59600.242836 5 29.29 29.21 28.32 28.02 2.46 0.17 2.610173
Osipovia 4986 652 246 59596.522535 5 27.46 27.37 25.76 24.92 2.56 0.21 5.611378
1991 EN 9857 1065 314 59598.310255 6 11.57 11.45 10.68 9.28 5.16 0.03 19.612052
@@
Table 1: Overview of the observations made by ATLAS and collected by rockAtlas, for a selected sample of 6 asteroids (this table is available in its entirety online as supplementary data and at this link). We provide the name and MPC number of each asteroid, alongside the number of detections in the $o$ and $c$-filters ($N_{o}$, $N_{c}$) and the date of the last observation used in this work, $\text{max}(t)$. The maximum phase angle observed in each filter is given by $\text{max}(\alpha_{o})$ and $\text{max}(\alpha_{c})$, and the maximum minus minimum phase angle range for each filter is $\Delta\alpha_{o}$, $\Delta\alpha_{c}$. Finally, the orbital elements obtained from astorb are also included; semimajor axis $a$, eccentricity $e$ and inclination $i$.
Name Number $n_{\textrm{app}}$ $t_{\textrm{app},0}$ (MJD) max($\alpha_{\textrm{app},o}$) max($\alpha_{\textrm{app},c}$) $N_{\textrm{fit},o}$ $N_{\textrm{fit},c}$ $H_{o}$ $H_{c}$ $\sigma_{Ho}$ $\sigma_{Hc}$ $G$ $\sigma_{G}$ Primary Apparition Keepo Keepc-o
Moguntia 766 0 57380.664873 17.90 18.00 47 59 9.28 9.66 6.93e-03 3.65e-03 0.21 NaN 0 1 1
Moguntia 766 1 57883.625009 18.09 17.34 146 20 10.09 10.67 8.86e-03 2.34e-02 0.21 NaN 0 1 0
Moguntia 766 2 58298.591274 20.52 20.53 199 49 9.25 9.61 2.00e-03 4.54e-03 0.21 NaN 0 0 0
Moguntia 766 3 58733.610938 21.24 21.22 232 70 10.15 10.51 8.96e-03 1.61e-02 0.21 NaN 0 1 1
Moguntia 766 4 59190.638252 18.29 18.46 244 138 9.27 9.63 2.42e-03 1.22e-03 0.21 2.87e-03 1 1 1
Priamus 884 0 57384.434340 9.92 9.94 41 74 8.59 8.99 1.93e-02 1.54e-02 0.23 NaN 0 0 0
Priamus 884 1 57673.634360 10.30 10.30 145 88 8.43 8.86 5.70e-03 1.13e-02 0.23 NaN 0 1 1
Priamus 884 2 58037.632049 10.94 10.78 179 22 8.39 8.72 3.90e-03 1.14e-02 0.23 NaN 0 0 0
Priamus 884 3 58460.627613 11.82 11.82 292 35 8.39 8.75 9.36e-03 9.19e-03 0.23 1.48e-02 1 1 1
Priamus 884 4 58867.638704 12.60 12.59 200 80 8.50 8.86 7.11e-03 1.27e-02 0.23 NaN 0 1 1
Priamus 884 5 59255.640729 12.05 12.37 20 14 8.55 8.96 2.54e-02 1.50e-02 0.23 NaN 0 0 0
Marietta 2144 0 57481.601935 20.26 20.39 45 112 10.99 11.39 2.12e-02 1.33e-02 0.21 NaN 0 1 1
Marietta 2144 1 57943.566586 21.55 21.36 175 27 11.06 11.39 1.14e-02 2.48e-02 0.21 NaN 0 1 0
Marietta 2144 2 58417.608911 20.14 20.28 256 52 10.96 11.42 8.33e-03 1.77e-02 0.21 NaN 0 1 1
Marietta 2144 3 58846.631123 19.45 19.45 309 126 11.08 11.41 8.92e-03 1.27e-02 0.21 NaN 0 1 1
Marietta 2144 4 59300.621863 20.58 20.47 410 90 11.02 11.41 1.28e-02 1.39e-02 0.21 1.58e-02 1 1 0
Yatskiv 2728 0 57400.538519 22.19 21.77 49 75 12.42 12.56 1.66e-02 2.09e-02 0.11 NaN 0 1 1
Yatskiv 2728 1 57826.587072 28.61 27.77 113 13 12.29 12.59 1.07e-02 1.65e-02 0.11 NaN 0 1 0
Yatskiv 2728 2 58385.616541 20.64 19.43 131 28 12.30 12.65 1.23e-02 3.43e-02 0.11 NaN 0 1 0
Yatskiv 2728 3 58790.634282 23.23 23.01 231 66 12.50 12.63 1.24e-02 1.82e-02 0.11 NaN 0 1 1
Yatskiv 2728 4 59276.616956 29.29 29.21 404 85 12.31 12.56 1.55e-02 1.09e-02 0.11 1.35e-02 1 1 1
Osipovia 4986 0 57591.583406 27.46 27.37 48 46 13.21 13.63 1.53e-02 1.65e-02 0.30 NaN 0 1 1
Osipovia 4986 1 58037.632049 24.71 23.86 158 37 13.23 13.63 1.01e-02 2.07e-02 0.30 NaN 0 1 1
Osipovia 4986 2 58515.672604 NaN NaN NaN NaN NaN NaN NaN NaN NaN NaN NaN 0 0
Osipovia 4986 3 58974.596053 25.95 25.79 191 96 13.19 13.58 1.82e-02 9.53e-03 0.30 2.31e-02 1 1 1
Osipovia 4986 4 59493.612303 26.30 24.83 110 35 13.22 13.60 1.10e-02 2.15e-02 0.30 NaN 0 1 0
1991 EN 9857 0 57461.642500 NaN NaN NaN NaN NaN NaN NaN NaN NaN NaN NaN 0 0
1991 EN 9857 1 57882.583924 9.72 8.05 36 7 10.04 10.48 2.03e-02 6.54e-02 0.16 NaN 0 0 0
1991 EN 9857 2 58269.545450 11.38 11.37 140 45 10.16 10.49 1.03e-02 1.41e-02 0.16 NaN 0 1 1
1991 EN 9857 3 58613.616782 11.57 11.45 255 66 10.09 10.40 4.74e-03 1.36e-02 0.16 NaN 0 1 1
1991 EN 9857 4 58997.589051 11.40 11.39 209 109 10.09 10.44 3.96e-03 7.08e-03 0.16 NaN 0 1 1
1991 EN 9857 5 59414.593958 11.17 10.99 352 46 10.10 10.46 1.36e-02 1.34e-02 0.16 2.14e-02 1 1 1
Table 2: Results of the phase curve fitting algorithm to determine phase curves for each apparition of the selected objects from table 1. NaN entries indicate where a phase curve fit could not be determined (e.g. the apparition was within the galactic plane and all observations were rejected). We assign an index to each apparition for an object ($n_{\text{app}}$), where the first observation in an apparition was made at time $t_{\text{app},0}$. The maximum phase angle observed in an apparition is given by $\text{max}(\alpha_{\text{app},o})$ and $\text{max}(\alpha_{\text{app},c})$ for the $o$ and $c$-filters respectively. The number of detections used to fit the phase curve (after outlier rejection) is denoted by $N_{\text{fit},o}$ and $N_{\text{fit},c}$. We show the fitted phase curve parameters for the $HG$ B89 system, alongside their uncertainties ($\sigma_{H}$, $\sigma_{G}$) in both $o$ and $c$-filters. The “Primary Apparition” flag indicates which apparition was fitted for both absolute magnitude and phase parameter. We also flag whether an apparition passed the quality control checks (sections 2.4 & 2.5) in the $o$ filter (“Keepo”) or if it passed in both $o$ and $c$ (“Keepc-o”). Only selected columns are displayed here. The table of apparitions is available in its entirety online as supplementary data and at this link, with additional columns describing in detail the phase curve fit to each apparition for both the B89 and P16 models.
@@
Figure 4: Log-scale distributions of $H$ (left), $G$ (middle) and $H_{c}-H_{o}$ colour (right) for ATLAS phase curve fits that have passed the quality control checks (section 2.5) and are included in the SAP dataset (section 3.1). $H$ and $G$ are given in the $o$-filter and B89 model. In both figures we show the underlying distributions of S and C type asteroids within the ATLAS dataset (taxonomy from SsODNet; Berthier et al., 2023). The median values of the S and C-type distributions are indicated by the vertical line of matching colour and their values are as follows: $H_{o}(\mathrm{S})=14.83$, $H_{o}(\mathrm{C})=14.52$, $G(\mathrm{S})=0.29$, $G(\mathrm{C})=0.21$, $H_{c}-H_{o}(\mathrm{S})=0.39$, $H_{c}-H_{o}(\mathrm{C})=0.27$. In the middle panel we indicate the boundaries on $G$ for a monotonically increasing B89 phase curve with vertical dashed lines.
@@
Figure 6: Two dimensional histogram of $H_{c}-H_{o}$ colour vs. $G$. The bimodal distribution of matched S and C-type asteroids (SsODNet; Berthier et al., 2023) is highlighted by the contoured regions, which indicates the 25% level with respect to the largest histogram bin value of their sub-distributions in $H_{c}-H_{o}$ vs. $G$ parameter space. The median values of the S and C-types with ATLAS colours are $G=0.29,0.21$ and $H_{c}-H_{o}=0.39,0.27\,$\mathrm{m}\mathrm{a}\mathrm{g}$$ respectively.
We have also derived the phase curve parameters of asteroids using the P16 $H_{12}G_{12}$ system, where we use the notation $H_{12}$ to denote absolute magnitudes in this system. As described in section 2.5 the quality control cuts were developed on the $HG$ dataset and then applied separately to the residuals of the $H_{12}G_{12}$ phase curve fits. Therefore not all asteroids are present in both datasets. Regardless, both phase curve models show broadly similar results. As before the phase parameter and colour distributions are both bimodal due to the dominant S and C-type asteroids, the former shown in figure 7. The quality control metrics developed for $HG$ do not lead to as strong restrictions on $G_{12}$ for “physical” values between 0 and 1. This appears to be mostly due to inherent differences between the $HG$ and $H_{12}G_{12}$ models. Figure 8 shows that there is a strong correlation between $G_{12}$ and $G$ for objects in our dataset with a phase curve fit in both models. We have verified that this is caused by the different formulations of the two models and is not due to systematic effects in our phase curve fitting procedure. As detailed in appendix B (figure 17), we generated a range of ideal model $HG$ phase curves with $0\leq G\ \leq 1$ and then fitted those simulated observations with the $H_{12}G_{12}$ model. The locus of best fit $G_{12}$ values for phase curves with parameter $G$ is shown in the top panel of figure 8. It can be seen that fitting P16 to observed phase curves naturally gives rise to a range of $G_{12}$ larger than $(0,1)$.
Figure 7: Log-scale distribution of P16 $G_{12}$ phase parameters for the quality controlled SAP dataset. Distributions of objects with S and C-type taxonomic classifications from Berthier et al. (2023) are also shown. Similar to figure 4 the median values of the S and C-type distributions are indicated by coloured vertical lines (with values of $G_{12}(\mathrm{S})=0.22$, $G_{12}(\mathrm{C})=0.51$) and the limits $10<G_{12}<$ 1 for a “physical” phase curve are shown as vertical dashed lines.
@@
Figure 8: Two dimensional histogram showing the relation between the $G_{12}$ and $G$ parameters for objects in our database (upper panel). A 3rd-order polynomial is fitted to the data as shown by the black curve (equation 9a). The circular markers indicate the values of $G_{12}$ when a $H_{12}G_{12}$ model is fitted to a $HG$ model (appendix B). This empirical test follows the trend in the ATLAS data closely. The bottom panel shows the two dimensional histogram of the difference in fitted absolute magnitude $H_{12}-H$ vs. $G$. Again, the curve is a 3rd-order polynomial fit (equation 9b) and the circular markers indicate the expected difference in absolute magnitude when $H_{12}G_{12}$ is fitted to $HG$. The horizontal line indicates the zero value of the $H_{12}-H$ axis.
Figure 9: Two dimensional histogram which shows the variation of $\Delta H=H_{12}-H$ with $H$. The median difference in absolute magnitude between the two phase curve models is $-0.11$ (indicated by the horizontal dashed line). There is a negative correlation between $\Delta H$ and $H$, as well as a trend for a larger dispersion in $\Delta H$ for fainter objects (see also figure 8). The horizontal solid line indicates the zero value of the $H_{12}-H$ axis.